# A Case Study of Cross-Lingual Zero-Shot Generalization for Classical Languages in LLMs

### v1

@@ Appendix A Question Answering Dataset
The broad semantic and domain-specific categories of the questions are detailed in Tables 4 and 4.
@@
Appendix B Retrieval Augmented Generation
@@
We evaluated RAG with different $k$ using GPT-4o with prompts in Sanskrit. The EM scores are plotted in Fig. 3. It can be seen that BM25 outperforms AvgFT and AvgGV, it may also be inferred from both plots that $k=4$ would be optimal. At this value of $k$, the performance of RAG in the case of Bhāvaprakāśanighaṇṭu already starts to deteriorate, although not significantly. Keeping both plots in mind, we have fixed the value of $k$ at 4.
The prompts for RAG are detailed in Appx. C.4.
Appendix C Prompts
The Sanskrit prompts are in Devanāgari script. In this appendix, we provide the prompts (transliterated in IAST scheme [^5] https://en.wikipedia.org/wiki/International_Alphabet_of_Sanskrit_Transliteration).
C.1 Prompts for Named Entity Recognition
Prompt in English
Recognize the named entities from the following sentence in {LANGUAGE}. The valid tags are {ENTITY TYPES}. Do not provide explanation and do not list out entries of ‘O’.
Example:{EXAMPLE}
(Note: EXAMPLE in Sanskrit is:
daśarathaḥ ayodhyāyāḥ rājā āsīt
{‘B-HUMAN’: [‘daśarathaḥ’], ‘B-CITY’: [‘ayodhyāyāḥ’]}
An example is provided to make the output structure clear.)
Prompt in Sanskrit
adho datta vākye named entities abhijānīhi (NER kuru). tadapi vivṛtam mā kuru, kevalam pṛṣṭa viṣayasya uttaram dehi.
entities eteṣu vargeṣu vartante - {ENTITY TYPES}
udāharaṇāya - {EXAMPLE}
C.2 Prompts for Machine Translation
Prompt in English
Prompt in Sanskrit
(Sanskrit QA Prompts)
In the following prompts, TOPIC is either ‘Rāmāyaṇa’ or ‘Āyurveda’.
C.3 Prompts for Closed-book QA
Prompt in English
{QUESTION} {CHOICES}
Prompt in Sanskrit
{QUESTION} {CHOICES}
C.4 Prompts for RAG-QA
Prompt in English
contexts: {CONTEXTS}
question:{QUESTION} {CHOICES}
Prompt in Sanskrit
sandarbhāḥ: {CONTEXTS}
praśnaḥ: {QUESTION} {CHOICES}
@@ Appendix E Supplementary Results
gpt-4o 0.51
llama-3.1-405b-instruct 0.49
NER (lat)
@@
In Table 5, we compare the out-of-domain performance of our evaluated models against previously reported fine-tuned models. For MT (san-eng), open-source model llama-3.1-405b-instruct achieves a BLEU score of 17.1, outperforming both Google Trans and IndicTrans, while for NER (lat), the performance of LLMs is roughly comparable to that of fine-tuned LatinBERT variants. Although fine-tuned models yield superior results on in-domain data, our findings indicate that multilingual LLMs are competitive—or even superior—in their zero-shot generalization.
@@ Appendix F LLMs with Knowledge Graphs
Arriving at an answer by an LLM integrated with a knowledge graph (KG) through Think-on-Graph (ToG) (Sun et al., 2024) paradigm involves several prompting steps for each hop from starting entity nodes. Firstly, the LLM lists entities from the input questions further lemmatized by our lemmatizer previously described. The relationships from and to these entities are then extracted by traversing the KG. The LLM then lists relationships with relevance scores, which are further used to prune the relationships, retaining only the best three. Unexplored entities connected by these relationships are then known from the KG, which are similarly pruned to retain the three most relevant ones. The LLM then reasons whether these extracted paths suffice to answer the given question. If no, the cycle is repeated, i.e., it traverses a hop further up to a depth $d$. Otherwise, the LLM answers using the context from the extracted paths.
@@ F.2 LLM-KG Prompts
api ca yathā’vaśyam adhaḥ dattaiḥ knowledge-graph-taḥ niṣkarṣita-viṣayaiḥ sahạyyaṃ gṛhāṇa. tattu sarvadā sādhu iti nā’sti pratītiḥ. uttaram yāvad laghu śakyam tāvat laghu bhavet.

### v2

@@ Appendix B Question Answering Dataset
The broad semantic and domain-specific categories of the questions are detailed in Tables 5 and 5.
@@
Appendix C Retrieval Augmented Generation
@@
The prompts for RAG are detailed in Appx. A.4.
@@ Appendix E Supplementary Results
gpt-4o 0.55
llama-3.1-405b-instruct 0.36
NER (lat) on Ars Amatoria dataset
@@
In Table 6, we compare the out-of-domain performance of our evaluated models against previously reported fine-tuned models. For MT (san-eng) on Mann ki Baat dataset (Maheshwari et al., 2024), open-source model llama-3.1-405b-instruct outperforms both Google Trans and IndicTrans, while for NER (lat) on Ovid’s Ars Amatoria dataset (Beersmans et al., 2023), the performance of gpt-4o is better than that of fine-tuned LatinBERT variants. Although fine-tuned models yield superior results on in-domain data, our findings indicate that multilingual LLMs are superior in their zero-shot generalization.
@@ Appendix F LLMs with Knowledge Graphs
Figure 5: Overview of augmenting a LLM with a knowledge graph (KG) through Think-on-Graph (ToG) paradigm.
Arriving at an answer by an LLM integrated with a knowledge graph (KG) through Think-on-Graph (ToG) (Sun et al., 2024) paradigm involves several prompting steps for each hop from starting entity nodes as illustrated in Fig. 5. Firstly, the LLM lists entities from the input questions further lemmatized by our lemmatizer previously described. The relationships from and to these entities are then extracted by traversing the KG. The LLM then lists relationships with relevance scores, which are further used to prune the relationships, retaining only the best three. Unexplored entities connected by these relationships are then known from the KG, which are similarly pruned to retain the three most relevant ones. The LLM then reasons whether these extracted paths suffice to answer the given question. If no, the cycle is repeated, i.e., it traverses a hop further up to a depth $d$. Otherwise, the LLM answers using the context from the extracted paths.
@@ F.2 LLM-KG Prompts
api ca yathā’vaśyam adhaḥ dattaiḥ knowledge-graph-taḥ niṣkarṣita-viṣayaiḥ sahāyyaṃ gṛhāṇa. tattu sarvadā sādhu iti nā’sti pratītiḥ. uttaram yāvad laghu śakyam tāvat laghu bhavet.
@@
F.3 LLM-KG Results
Method gpt-4o claude-3.5-sonnet gemini-1.5-pro mistral-large-2 llama-3.1-405b-instruct
Closed-book 0.381 0.242 0.148 0.333 0.346
RAG-BM25 0.478 0.521 0.459 0.434 0.323
LLM-KG 0.381 0.254 - 0.341 -
Table 7: Exact Match (Scores) of various models (including those not part of main experiments) in Sanskrit Question-Answering task (Sanskrit Prompts) with LLM-KG paradigm compared against zero-shot and RAG-BM25 paradigms.
The LLM-KG paradigm was evaluated exclusively using Sanskrit prompts on the two QA datasets and included additional models not part of the main experiments—namely, claude-3.5-sonnet (AnthropicAI, 2024), gemini-1.5-pro (Google, 2024), and mistral-large-2 (MistralAI, 2024). Table 7 presents the results in comparison with the closed-book and RAG-BM25 paradigms. Overall, performance gains from closed-book to LLM-KG are modest and fall short of the improvements observed with RAG. This may be partly attributed to the complexity of the LLM-KG setup, which requires multi-step prompting and adherence to a structured output format. Notably, models like gemini-1.5-pro and llama-3.1 frequently fail to follow this structured format, rendering them ineffective for running ToG. The strict formatting requirements may also pose challenges for other models, particularly those less adapted to Sanskrit. Interestingly, while claude-3.5-sonnet achieves the best results with RAG-BM25, it lags behind gpt-4o and mistral-large-2 in both the closed-book and LLM-KG paradigms.
Method gpt-4o claude-3.5-sonnet mistral-large-2
closed-book 0.32 0.21 0.25
LLM-KG 0.34 0.34 0.35
Method gpt-4o claude-3.5-sonnet mistral-large-2
closed-book 0.40 0.25 0.36
LLM-KG 0.39 0.23 0.34
Table 8: Comparison of Exact Match (EM) scores between closed-book and LLM-KG paradigms for selected questions when the answer (a) can likely be inferred from KG and (b) cannot be inferred from KG.
Table 8 presents a breakdown of performance based on whether the question topics are covered in the current KG—specifically, the kingdoms category (27 questions) in the Rāmāyaṇa dataset and the annotated chapters (299 questions) in Bhāvaprakāśanighaṇṭu. For these subsets, which are likely answerable from the KG, LLM-KG shows clear improvements over the closed-book setting, indicating that access to a near-complete KG can significantly enhance performance. In contrast, for questions outside these categories or chapters, no such improvement is observed, reinforcing the hypothesis that KG completeness is crucial for the effectiveness of LLM-KG. Determining domains where knowledge graphs may outperform or be more appropriate than RAG remains an open question for future research.
Appendix G Categories for Named Entity Recognition
The categories for NER in Sanskrit, Ancient Greek, and Latin, along with their rough translation and brief explanations, wherever applicable, are provided here.
Entity Type Translation Description
Manuṣya Human A mortal human being
Deva Deity Divine celestial being; god or goddess
Gandharva $\sim$ Heavenly musician in the service of the gods
Apsaras $\sim$ Beautiful female spirits known for dance and charm
Yakṣa $\sim$ Guardian spirit of natural treasures.
Kinnara $\sim$ Certain Semi-divine beings
Rākṣasa $\sim$ Malevolent being
Asura Anti-god Powerful beings opposed to the gods
Vānara Monkey-being Monkey-like humanoid
Bhallūka Bear-being Bear or Bear-like humanoid
Gṛdhra Vulture-being Vulture-like being
Ṛkṣa Bear-being Bear-like humanoid
Garuḍa Eagle-being Eagle-like being
Nāga Serpent-being Semi-divine serpent race
Svarga Heaven Abode of the gods
Naraka Hell Realm of punishment after death
Nadī River Flowing body of freshwater
Sāgara Sea Vast saltwater body
Sarovara Lake Large inland water body
Kūpa Well Man-made water source
Tīra Riverbank Edge or shore of a river
Dvīpa Island Land surrounded by water
Parvata Mountain Large natural elevation of earth
Nagara City Urban settlement or metropolis
Tīrtha Sacred Place Holy pilgrimage spot, often near water
Grāma Village Small rural settlement
Rājya Kingdom Territory ruled by a king
Vana Forest Dense growth of trees; wilderness
Udyāna Garden Cultivated green space
Marubhūmi Desert Dry, arid region
Prāsāda Palace Royal residence
Mandira Temple Sacred structure for worship
Āśrama Hermitage Secluded place for spiritual practice
Gṛha House Dwelling or home
Kuṭīra Hut Small and simple shelter
Guhā Cave Natural underground chamber
Mārga Road Pathway or route
Ratha Chariot Two- or four-wheeled ancient vehicle
Vimāna Airborne Vehicle Flying chariot or aircraft
Khadga Sword Bladed weapon
Dhanus Bow Weapon for shooting arrows
Bāṇa Arrow Projectile shot from a bow
Cakra Discus Spinning circular weapon
Gadā Mace Blunt weapon, often spiked
Tomara Javelin Thrown spear or missile
Śūla Spear Long-shafted piercing weapon
Kavaca Shield Defensive armor piece
Kañcuka Armor Protective body gear
Paraśu Axe Bladed tool/weapon
Astra Divine Weapon Supernatural weapon, often invoked
Ābharaṇa Ornament Decorative jewelry
Śaṅkha Conch Sacred spiral shell
Vādya Musical Instrument Instrument used in music
Nāṇa Currency Form of money or coin
Kula Clan Extended family or lineage
Jāti Species Species/Socio-economical Group
Gaṇa Tribe / Group Assembly or community
Ṛtu Season Climatic period of the year
Saṃvatsara Year Vedic year cycle
Māsa Month Lunar or solar month
Tithi Lunar Day Phase in the moon’s waxing/waning
Pakṣa Fortnight Half of a lunar month
Ayana Solstice Cycle Six-month movement of the sun
Yuga Epoch Cosmic age or era
Yoga Astronomical Combination Planetary conjunction
Karaṇa Half of Tithi Subdivision of a lunar day
Muhūrta Moment / Auspicious Time Small unit of time (about 48 minutes)
Lagna Ascendant Zodiac rising at time of birth
Graha Planet Celestial influencer
Nakṣatra Lunar Mansion One of 27 lunar constellations
Rāśi Zodiac Sign Segment of the zodiac
Dhuma-ketu Comet Celestial object with a tail
Utsava Festival Celebratory event
Pūjā Worship Ritual offering and prayer
Yajña Vedic Sacrifice Sacred fire ritual
Upacāra Ritual Offering Ceremonial gesture or item
Saṃskāra Life-Cycle Rite Hindu ritual of life transition
Aniścita Undecided Something that is not yet determined
Vṛkṣa Tree Large woody plant
Guccha Shrub Small bushy plant
Lata Vine Climbing or trailing plant
Puṣpa Flower Blossom of a plant
Phala Fruit Edible plant product
Patra Leaf Green foliage part
Stambha Stem Main structural plant part
Tvak Bark Outer layer of tree
Mūla Root Underground part of plant
Pakṣī Bird Feathered flying animal
Sarpa Snake Legless reptile
Table 9: Entity types occuring in Sanskrit NER
Entity Type Description
NORP Ethnic groups, demonyms, schools
ORG Organizations
GOD Supernatural beings
LANGUAGE Languages and dialects
LOC Cities, empires, rivers, mountains, and so forth.
PERSON Individual persons
Table 10: Entity types occuring in Ancient Greek NER (Myerston, 2025). The types without descriptions—EVENT and WORK—have very few occurances in the dataset.
Entity Type Description
PER Person
LOC Locations, places
GRP Other groups such as tribes
Table 11: Entity types occuring in Latin NER are quite standard types.